Shift summary for day, July 14,99

Angela Bellavance (bellavan@rice.edu)
Wed, 14 Jul 1999 15:36:53 -0500 (CDT)

Shift Summary for day shift, 14/07/99

On Shift: N. Solomy, A. Bellavance

Hours of beam delivered at all intensities: 3:41
Hours of beam delivered at nominal intensity: 2:10
Hours of E832 data (special runs not included): 2:49
Hours of special runs (laser scans, muon runs...): 0:25

Accounting of lost time:

-Bad beam on startup
-Muon run on startup
-Access for BA, HA repair
-Problems with key tree loop "A"

========================================================================
Runs ended this shift:

Run 13902 - 535k events - COSMICS
Run 13903 - 1.7M events - BEAM_MUON
Run 13904 - 1.45M events - BEAM_832
Run 13905 - aborted
Run 13906 - Ped. run
Run 13907 - ongoing - BEAM_832

Tapes brought to FCC this shift: none
(including tapes for online split)
========================================================================
Accesses during shift:

RickK and Elizabeth:
- BASUM: put back old BASUM ss to get -20mV DC offset
- HA Sigma: DC offset was +100mV! Adjusted to zero via front
panel screws on 428s.
- HA Sigma: moved from "DC coupled" ADC to HA(29) ADC cable.

========================================================================
Overview of shift:
Wait for return of beam
E832 datataking
access for BA,HA
more E832 datataking

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
13902 7:00 Message still on screen from before: "Beam fallout
in MI. No estimate."
8:26 TV (ch. 13) starts making obnoxious noises.
13902 8:54 MCR calls. "Beam on it's way." Stop cosmic run,
setup for muon run.
9:15 Quick search under floor tiles and above ceiling
tiles for source of bad smell turns up nothing.
9:39 Message: "Having difficulty establishing beam.
Experts investigating. No estimate."
13903 1 9:47 Start muon run.
9:53 Bry air in the furnace room is noticed as being
turned on. Turned off, should stay off until
you expect HV back on TRDs.
13903 26 10:18 Beam finally shows up.
13903 10:40 Stop muon run. Ramping magnets for data taking.
Ran regmon#muonrun. Minor problems refered to
Elizabeth.
13904 1 11:00 Beam_832 run started.
13904 11:05 Forgot to go a KFREND; did it now.
13904 8 11:08 Forgot to turn triggers back on after KFREND; done
now.
13904 11:51 Stop run in preparation for access.
12:00 RickK and Elizabeth go on access for the BA and
HA.
13906 1 12:33 Ped. run started. Come out of access.
13906 12:36 Ped. run stopped. Peds updated.
13907 1 12:47 Run started but problems with key tree.
13907 11 13:00 Key Tree fixed, beam returns.
13:32 all checklists completed. In doing the run
checklist, CsI index 961 had a comparitor problem
(freq:9; range:3; cap:2) but the live-time still
looks fine. PeterSh and Ed notified.
13:58 Index 961 freq. now 11.
14:30 Index 961 freq. now 18.
13907 92-94 DYC 3-53 full. KFREND fixes.
13907 103- 15:06- SY energy conservation.
110 15:15
13907 113-? 15:19-? SY energy conservation.
13907 15:25 Message: "Accelerator problems. Holding off
intensity. Est. 15 min."
13907 15:30 Hand off to eve shift, still no beam.

========================================================================
Follow up Notes on Previous Shift:

-Muon run done on start-up.
-SDAQ red buttons "RR160" and "UCLA-PS4" can be ignored for now.

========================================================================
Items to Follow up on Next Shift:

-Key Tree problem the problem was that "Loop A" on the key tree would not
close/light. If a problem in future, there is a depressor switch to
activate this on the left inside edge of the box (door open). The top of
this switch is a screw, which sometimes gets turned in too far. Turn it
CCW to raise it back out and try closing the door again.

-SDAQ red buttons "RR160" and "UCLA-PS4" can be ignored for now.

****NOTE TO ALL: When writing down the magnet polarity in the run
logbook, make sure you take it directly off the Epicure readout screen.

_______
Angela Bellavance work: (630) 840-4144
bellavan@rice.edu home: (630) 548-3412
http://www.owlnet.rice.edu/~bellavan/ page: (630) 722-6861
"It has yet to be proven that intelligence has any survival value."
- Arthur C. Clarke